# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

### v1

@@ 2.2.1 The number of particles near maximum
where in the last step we use $0\leqslant c_{2}-\lambda_{n}\ll(\log n)/n$. In a similar manner, (while using (6) of [13]), we obtain the estimate
@@
×$\hskip 113.81102pt\times\left(jxe^{c_{2}(x-j)+3((s\log n)/n-\log s)/2}\right)^{2}$
@@
where in the last step we argue similarly as in the proof of Lemma 2.7 of [13]. Combined with (16), (17), and the Paley-Zygmund inequality leads to (13). ∎
@@ 2.2.2 Large deviation computations
In the following, we let $\{(X_{j},\mathbf{Y}_{j})\}_{j\in\mathbb{N}}$ be a sequence of i.i.d. random variables with the same distribution as ${\bm{\xi}}$, and $(X,\mathbf{Y})=\sum_{j=1}^{n}(X_{j},\mathbf{Y}_{j})$. Recall the definition of $m_{n}$ from (8).
@@
    $\mathbb{P}(\mathbf{Y}\in B_{\mathbf{0}}(u(n))\mid X\geqslant m_{n}+c)\ll u(n)^{d-1}n^{-(d-1)/2}.$ (20)
@@
    Fix $a(n)\to\infty$ and $a(n)=o(\log n)$. Uniformly for $\mathbf{y}$ with $\left\lVert\mathbf{y}\right\rVert\leqslant C\sqrt{n}$ and $c\in[-C\log n,C\log n]$,
    $\mathbb{P}(\mathbf{Y}\in B_{\mathbf{y}}(1)\mid X\in[m_{n}+c,m_{n}+c+a(n)])\gg n^{-(d-1)/2}.$ (21)
@@
    $\mathbb{P}(\mathbf{Y}\not\in B_{\mathbf{0}}(K\sqrt{n})\mid X\geqslant m_{n}+c)<\varepsilon.$ (22)
@@ 2.2.3 Proof of Theorem 1 conditioned on Theorem 2
In view of Theorem 2, it suffices to prove that for any $a(x)$ that tends to infinity, and any $\varepsilon>0$, there exists $N>0$ such that for any $x\geqslant N$,
@@
(#$\mathbb{P}\big{(}\#W_{(\log x)^{2}}<L_{1}x^{({d-1})/{2}}\big{)}<1-\frac{1}{L}.$ (25)
Indeed, let
$\widetilde{W}_{n}:=\left\{v\in V_{n}:\eta_{v,n}(n)\in\Big{(}m_{n}-\alpha(x),m_{n}\Big{)}\right\}.$
It follows from Proposition 8 that for $x$ large enough, (#$\mathbb{P}(\#\widetilde{W}_{n}<L_{1}x^{(d-1)/2})<1-1/L$. Moreover, a union bound gives that out of $L_{1}x^{(d-1)/2}$ particles in $\widetilde{W}_{n}$, the probability that the last $d-1$ coordinates fall outside $B_{\mathbf{0}}(\sqrt{x})$ for some of them is at most $L_{1}x^{(d-1)/2}\mathbb{P}(\left\lVert\bf Z\right\rVert\geqslant\sqrt{x}/\log x)=o(1)$, where $\bf Z$ denotes a standard Gaussian random variable on $\mathbb{R}^{d-1}$. This proves (25).
@@
The plan is to evolve (on the event of (25)) the $L_{1}x^{(d-1)/2}$ particles in $W_{(\log x)^{2}}$ (independently) so that at least one of those families has a descendant landing in $B_{x}$ at time $A_{1}(x)+a(x)$ with high probability, where we recall that $B_{x}$ is the ball of radius one centered at $(x,\mathbf{0})$. Let us now fix $w\in W_{(\log x)^{2}}$. Define $\tau^{\prime}(x):=A_{1}(x)+a(x)-(\log x)^{2}$ and
$p(w):=\mathbb{P}\left(\bm{\eta}_{v,\tau^{\prime}(x)}(\tau^{\prime}(x))\in B_{x}-\bm{\eta}_{w,(\log x)^{2}}((\log x)^{2})\text{ for some }v\in V_{A_{1}(x)+a(x)-(\log x)^{2}}\right),$
@@
which is the probability that at least one descendant of $w$ lands in $B_{x}$ at time $A_{1}(x)+a(x)$. Here and later, we will without loss of generality assume that quantities that tend to infinity are integers. By independence and by picking the constant $L_{1}$ in (25) large enough, it suffices to prove $p(w)\gg x^{-(d-1)/2}$ for $w\in W_{(\log x)^{2}}$. Since $\eta_{w,(\log x)^{2}}((\log x)^{2})\in(m_{(\log x)^{2}}-\alpha(x),m_{(\log x)^{2}})$ and $m_{\tau^{\prime}(x)}\geqslant x-(m_{(\log x)^{2}}-\alpha(x))$, by the main result of [39],
@@
$\mathbb{P}\left(\eta_{v,\tau^{\prime}(x)}(\tau^{\prime}(x))\in B_{x}(1)-\eta_{w,(\log x)^{2}}((\log x)^{2})\text{ for some }v\in V_{A_{1}(x)+a(x)-(\log x)^{2}}\right)\geqslant\frac{1}{L}.$
Conditionally on existence, applying Proposition 10(ii) to the particle $v\in V_{A_{1}(x)+a(x)-(\log x)^{2}}$ with $\eta_{v,\tau^{\prime}(x)}(\tau^{\prime}(x))\in B_{x}(1)-\eta_{w,(\log x)^{2}}((\log x)^{2})$ leads to $p(w)\geqslant x^{-(d-1)/2}/L$. This establishes (23).
@@
Proof of the lower bound. Our goal is to bound $\mathbb{P}(\tau_{x}\leqslant A_{2}(x)-a(x))$ from above. According to Propositions 8 and 10, at time $A_{2}(x)$ there will be less than $O(1)$ number of particles that lie in $[x,\infty)\times B_{\mathbf{0}}(1)$ in expectation. We need to argue that other particles that stayed in $B_{x}$ earlier but currently do not stay in $[x,\infty)\times B_{\mathbf{0}}(1)$ do not contributed asymptotically. This is because most of the $O(1)$ particles will arrive not much earlier than $m_{1}(x)+m_{2}(x)$, and will not move quickly.
@@
Consider $n=A_{2}(x)$. Recall that $\varepsilon>0$ and $a(x)\to\infty$ are fixed. We define $T_{v,n}(x):=\min\{0\leqslant k\leqslant n:\eta_{v,n}(k)\geqslant x\}$ and for an integer $j\in[a(x),m_{2}(x)]$,
@@
$\leqslant\mathbb{P}\left(\bigcup_{j=a(x)}^{m_{2}(x)+L_{2}}\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x}\}\right)+\mathbb{P}\left(M_{A_{2}(x)-m_{2}(x)-L_{2}}\geqslant x\right)$
$\leqslant\sum_{j=a(x)}^{m_{2}(x)+L_{2}}\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x}\}\right)+\frac{\varepsilon}{4}.$ (26)
Let us further decompose the event $\{\bm{\eta}_{v,n}(k)\in B_{x}\}$ depending on the location at the first time the random walk $\{\bm{\eta}_{v,n}(k)\}_{0\leqslant k\leqslant n}$ reaches $[x,\infty)\times\mathbb{R}^{d-1}$, whether $\bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)$ with some $c_{3}>0$ to be determined.
@@
$\hskip 14.22636pt\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x},\leavevmode\nobreak\ \bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right)$
@@
(#$\mathbb{P}\left(\#\{v\in V_{A_{2}(x)-j}:\eta_{v,A_{2}(x)-j}(A_{2}(x)-j)\geqslant x\}\geqslant L_{3}j^{2}e^{-c_{2}c_{1}j}n^{(d-1)/2}\right)<\frac{\varepsilon}{10j^{2}}.$
@@
$\mathbb{P}\left(\bigcup_{v\in V_{A_{2}(x)-j}}\{\eta_{v,A_{2}(x)-j}(A_{2}(x)-j)\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right)\leqslant\frac{\varepsilon}{10j^{2}}+L_{3}j^{2}e^{-c_{2}c_{1}j}(c_{3}j)^{d-1}.$
Second, using that $j\geqslant a(x)$ and $a(x)\to\infty$ as $x\to\infty$, as well as Theorem 3.2 of [29], we have for some $c_{3},c_{4}>0$ large and independent of $x$, such that for all $x$ large enough,
$\hskip 14.22636pt\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x},\leavevmode\nobreak\ \bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)^{c}\}\right)$
@@
$\leqslant d\,\mathbb{P}\left(\max_{v\in V_{j}}|\eta_{v,j}(j)|\geqslant\frac{(c_{3}-1)j}{\sqrt{d}}\right)$
@@
For any $\varepsilon>0$ and $a(x)\to\infty$, the right-hand side is $<\varepsilon$ for $x$ large enough. This completes the proof of the lower bound (24).
@@ 2.3 Tightness of first passage times: proof of Theorem 2
We build upon ideas from [41], while in contrast, we do not assume uniform boundedness of the increments. [^6] In fact, we expect that our approach also leads to a concentration result for the maximum of BRW without assuming uniform boundedness of the increments. In the following, we write $C_{1},C_{2},C_{3},\dots$ as large constants and $c_{3},c_{4},\dots$ as small constants that may depend on the law of ${\bm{\xi}}$ and the branching rate. [^7] Not to confuse with the fixed constants $c_{1},c_{2}$ defined in Section 1.1.1. Here, we do not assume that ${\bm{\xi}}$ is spherically symmetric. We start with a few preparatory lemmas.
@@
Assume (A6). For any $C_{1}>0$ and $c_{3}>0$, there exists $C_{2}>0$ such that the following holds for $k$ large enough: let $(X_{j},\mathbf{Y}_{j})$ be a sequence of i.i.d. random variables with the same distribution as ${\bm{\xi}}$, $C>C_{2}$, and $(X,\mathbf{Y})=\sum_{j=1}^{Ck}(X_{j},\mathbf{Y}_{j})$. Then uniformly for $k>0$ and ${\mathbf{z}}\in\mathbb{R}^{d}$ such that $\left\lVert{\mathbf{z}}\right\rVert\leqslant C_{1}k$,
$\mathbb{P}((X,\mathbf{Y})\in B_{{\mathbf{z}}})\gg e^{-c_{3}k}.$
@@
(#$\mathbb{P}(\#\{v\in V_{n}:\left\lVert\bm{\eta}_{v,n}(n)\right\rVert\leqslant c_{5}n\}<e^{c_{6}n}\mid S)\ll e^{-c_{7}n}.$ (27)
Indeed, given $c_{5}$, let $c_{8}=c_{5}/(6d\bar{c}_{1})$ where here $\bar{c}_{1}$ stands for the maximum linear speed of the BRW across all $d$ directions, so that by Theorem 3.2 of [29], [^8] Theorem 3.2 of [29] is stated in the Schröder case, where $p_{0}+p_{1}>0$. However, the proof for the upper bound of the large deviation probability does not rely on this assumption; see Remark 3.3(b) therein.
$\mathbb{P}\left(\exists v\in V_{c_{8}n},\leavevmode\nobreak\ \left\lVert\bm{\eta}_{v,c_{8}n}(c_{8}n)\right\rVert>\frac{c_{5}n}{2}\right)\ll e^{-c_{7}n}.$ (28)
Conditioning upon the survival event $S$, with an exponentially small probability we have that #$\#V_{c_{8}n}\geqslant e^{c_{9}n}$. Evidently, each $v\in V_{c_{8}n}$ will lead independently with probability $>1/2$ to a particle $w\in V_{n}$ where $\left\lVert\bm{\eta}_{w,n}(n)-\bm{\eta}_{w,n}(c_{8}n)\right\rVert\leqslant c_{5}n/2$. This along with standard binomial estimates proves (27).
[…]
@@
Note that such a fixed path exists for at least half of $v_{j}$’s with overwhelming probability. In conclusion, we evolved at least $e^{c_{6}k}/2$ many particles independently at time $k$ for time $C_{2}k$, each resulting in a probability of $\gg e^{-c_{6}k/2}$ to land in $[k,\infty)\times\mathbb{R}^{d-1}$. A standard binomial estimate yields that with overwhelming probability, there exist $e^{c_{6}k/4}$ particles at time $(C_{2}+1)k$ present in $[k,\infty)\times\mathbb{R}^{d-1}$ (conditional on survival). This completes the proof. ∎
@@
In the proof of Theorem 2, we wish to use Lemma 13 to bound probabilities of the form $\mathbb{P}(\tau_{x}\geqslant\alpha,\leavevmode\nobreak\ \tau_{y}\leqslant\beta)$, where $\beta<\alpha$. This is intuitively possible since we may run the particle $v_{y}$ that reaches $B_{y}$ and bound the probability that it never reaches $B_{x}$ in a time longer than $\alpha-\beta$, where we have used the strong Markov property of the random walk. There is, however, a subtle issue for the survival of the branching process initiated by the particle $v_{y}$, as this event is almost independent of the survival event $S$ we are conditioning on (in the proof of Theorem 2). We will bypass this difficulty by excluding this termination event a priori, or by first having an exponential number $e^{cn}$ of particles in $B_{y}$ instead of a single particle $v_{y}$.
@@
Let $k=y/C_{3}$ and set $C_{5}\gg a+C_{6}+C_{7}$, all to be determined. Denote by $\tau^{(c_{11})}_{x+k}$ the waiting time until $e^{c_{11}k}$ many particles reach the ball $B_{x+k}$. Now note that
$\begin{split}\mathbb{P}(\tau_{x}>s-1)&\leqslant\mathbb{P}(s-1<\tau_{x}<s+C_{5}k)+\mathbb{P}\left(\tau^{(c_{11})}_{x+k}>s+C_{3}k+1+C_{6}C_{7}k\right)\\ &\hskip 71.13188pt+\mathbb{P}\left(\tau_{x}\geqslant s+C_{5}k,\leavevmode\nobreak\ \tau^{(c_{11})}_{x+k}\leqslant s+C_{3}k+1+C_{6}C_{7}k\right).\end{split}$ (31)
@@
(#$\mathbb{P}(\#\{v\in V_{C_{3}k}:\eta_{v,C_{3}k}(C_{3}k)\geqslant k\}<e^{c_{4}k})\leqslant q+o(e^{-c_{4}k}).$

### v2

@@ 2.2.1 The number of particles near maximum
where in the last step we use $0\leqslant c_{2}-\lambda_{n}\ll(\log n)/n$. In a similar manner, (while using (6) of [bramson2016convergence]), we obtain the estimate
@@
$\hskip 113.81102pt\times\left(jxe^{c_{2}(x-j)+3((s\log n)/n-\log s)/2}\right)^{2}$
@@
where in the last step we argue similarly as in the proof of Lemma 2.7 of [bramson2016convergence]. Combined with (16), (17), and the Paley-Zygmund inequality leads to (13). ∎
@@ 2.2.2 Large deviation computations
In the following, we let $\{(X_{j},\mathbf{Y}_{j})\}_{j\in\mathbb{N}}$ be a sequence of i.i.d. random variables with the same distribution as $\bm{\xi}$}, and $(F_{n},\mathbf{J}_{n})=\sum_{j=1}^{n}(X_{j},\mathbf{Y}_{j})$. Recall the definition of $m_{n}$ from (8). We may slightly abuse notation and use $B_{\mathbf{x}}(r)$ to denote the ball centered at $\mathbf{x}\in\mathbb{R}^{d-1}$ with radius $r$ in dimension d1d-1 (instead of dimension $d$).
@@
    $\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\mid F_{n}\geqslant m_{n}+c)\ll u(n)^{d-1}n^{-(d-1)/2}.$ (20)
@@
    Fix $a(n)\gg 1$ and $a(n)=o(\log n)$. Uniformly for $\mathbf{y}$ with $\left\lVert\mathbf{y}\right\rVert\leqslant C\sqrt{n}$ and $c\in[-C\log n,C\log n]$,
    $\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1)\mid F_{n}\in[m_{n}+c,m_{n}+c+a(n)])\gg n^{-(d-1)/2}.$ (21)
@@
    $\mathbb{P}(\mathbf{J}_{n}\not\in B_{\mathbf{0}}(K\sqrt{n})\mid F_{n}\geqslant m_{n}+c)<\varepsilon.$ (22)
@@ 2.2.3 Proof of Theorem 1 conditioned on Theorem 2
In view of Theorem 2, it suffices to prove that there exists $L>0$ such that for any $a(x)$ that tends to infinity, and any $\varepsilon>0$, there exists $N>0$ such that for any $x\geqslant N$,
@@
$\mathbb{P}\big(\#W_{(\log x)^{2}}<L_{1}x^{({d-1})/{2}}\big)<1-\frac{1}{L},$ (25)
for some constant $L>0$ depending on $L_{1}$ only. Indeed, let
$\widetilde{W}_{n}:=\left\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\in\Big(m_{n}-\alpha(x),m_{n}\Big)\right\}.$
It follows from Proposition 8 that for $x$ large enough, with $n=(\log x)^{2}$, $\mathbb{P}(\#\widetilde{W}_{n}<L_{1}x^{(d-1)/2})<1-1/L$. Comparing the definitions of $W_{n}$ and $\widetilde{W}_{n}$, it remains to show that at generation $n=(\log x)^{2}$, the maximum norm of the BRW is smaller than $\sqrt{x}$ with probability 1o(1)1-o(1). But this is a direct consequence of Theorem 1.1 of [zhang2024large].
@@
The plan is to evolve (on the event of (25)) the $L_{1}x^{(d-1)/2}$ particles in $W_{(\log x)^{2}}$ (independently) so that at least one of those families has a descendant landing in $B_{x}$ at time $A_{1}(x)+a(x)$ with high probability, where we recall that $B_{x}$ is the ball of radius one centered at $(x,\mathbf{0})$. Let us now fix $w\in W_{(\log x)^{2}}$. Define $\tau^{\prime}(x):=A_{1}(x)+a(x)-(\log x)^{2}+K_{1}$ for some large constant $K_{1}>0$ (to be determined) and
$p(w):=\mathbb{P}\left(\bm{\eta}_{v,\tau^{\prime}(x)}(\tau^{\prime}(x))\in B_{x}-\bm{\eta}_{w,(\log x)^{2}}((\log x)^{2})\text{ for some }v\in V_{\tau^{\prime}(x)}\right),$
@@
which is the probability that at least one descendant of $w$ lands in $B_{x}$ at time $A_{1}(x)+a(x)+K_{1}$, or equivalently, the probability that a BRW reaching the set $B_{x}-\bm{\eta}_{w,(\log x)^{2}}((\log x)^{2})$ at generation $\tau^{\prime}(x)$. By independence and by picking the constant $L_{1}$ in (25) large enough, it suffices to prove $p(w)\gg x^{-(d-1)/2}$ for $w\in W_{(\log x)^{2}}$. Note that $\eta^{(1)}_{w,(\log x)^{2}}((\log x)^{2})\in(m_{(\log x)^{2}}-\alpha(x),m_{(\log x)^{2}})$ and that for $K_{1}$ chosen large enough,
$m_{\tau^{\prime}(x)}$ $=x+\frac{d-1}{2c_{2}}\log x+\frac{2}{c_{2}}\log\log x-c_{1}(\log x)^{2}+c_{1}a(x)+c_{1}K_{1}+\frac{3\log c_{1}}{2c_{2}}+o(1)$
$\geqslant x+\frac{d-1}{2c_{2}}\log x+\frac{2}{c_{2}}\log\log x-c_{1}(\log x)^{2}+c_{1}a(x)+1$
$=x-(m_{(\log x)^{2}}-\alpha(x))+1.$
Therefore, the asymptotic of the maximum at time $\tau^{\prime}(x)$, $m_{\tau^{\prime}(x)}$, is at least $x+1-\eta^{(1)}_{w,(\log x)^{2}}((\log x)^{2})$. Using the lower bound on $\Delta_{n,x}$ derived in the proof of (13), there is some constant $L>0$ such that
@@
$\mathbb{P}\left(\eta^{(1)}_{v,\tau^{\prime}(x)}(\tau^{\prime}(x))\in\big(x-\frac{1}{2},x+\frac{1}{2}\big)-\eta^{(1)}_{w,(\log x)^{2}}((\log x)^{2})\text{ for some }v\in V_{\tau^{\prime}(x)}\right)\geqslant\frac{1}{L}.$
Conditioned on the above event, Proposition 10(ii) shows that the particle $v$ that realizes this event has a chance of $\gg x^{-(d-1)/2}$ to land in $B_{x}-\bm{\eta}_{w,(\log x)^{2}}((\log x)^{2})$ at time $\tau^{\prime}(x)$, leading to $p(w)\geqslant x^{-(d-1)/2}/L$. This establishes (23).
@@
Proof of the lower bound. Our goal is to bound $\mathbb{P}(\tau_{x}\leqslant A_{2}(x)-a(x))$ from above. According to Propositions 8 and 10, at time $A_{2}(x)$ there will be less than $O(1)$ number of particles that lie in $x,\infty)\times B_{\mathbf{0}}(1)$ in expectation. We need to argue that other particles that stayed in $B_{x}$ earlier but currently do not stay in $x,\infty)\times B_{\mathbf{0}}(1)$ do not contribute asymptotically. This is because most of the $O(1)$ particles will arrive not much earlier than $t_{1}(x)+t_{2}(x)$, and will not move quickly.
@@
Consider $n=A_{2}(x)$. Recall that $\varepsilon>0$ and $a(x)\to\infty$ are fixed. We define $T_{v,n}(x):=\min\{0\leqslant k\leqslant n:\eta^{(1)}_{v,n}(k)\geqslant x\}$ and for an integer $j\in[a(x),t_{2}(x)]$,
@@
$\leqslant\mathbb{P}\left(\bigcup_{j=a(x)}^{t_{2}(x)+L_{2}}\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x}\}\right)+\mathbb{P}\left(M_{A_{2}(x)-t_{2}(x)-L_{2}}\geqslant x\right)$
$\leqslant\sum_{j=a(x)}^{t_{2}(x)+L_{2}}\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x}\}\right)+\frac{\varepsilon}{4}.$ (26)
Let us further decompose the event $\{\bm{\eta}_{v,n}(k)\in B_{x}\}$ depending on the location at the first time the random walk $\{\bm{\eta}_{v,n}(k)\}_{0\leqslant k\leqslant n}$ reaches $x,\infty)\times\mathbb{R}^{d-1}$, whether $\bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)$ with some $c_{3}>0$ to be determined.
@@
$\hskip 14.22636pt\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x},\penalty 10000\ \bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right)$
@@
$\mathbb{P}\left(\#\{v\in V_{A_{2}(x)-j}:\eta^{(1)}_{v,A_{2}(x)-j}(A_{2}(x)-j)\geqslant x\}\geqslant L_{3}j^{2}e^{-c_{2}c_{1}j}n^{(d-1)/2}\right)<\frac{\varepsilon}{10j^{2}}.$
@@
$\mathbb{P}\left(\bigcup_{v\in V_{A_{2}(x)-j}}\{\eta^{(1)}_{v,A_{2}(x)-j}(A_{2}(x)-j)\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right)$
$\hskip 170.71652pt\leqslant\frac{\varepsilon}{10j^{2}}+L_{3}j^{2}e^{-c_{2}c_{1}j}(c_{3}j)^{d-1}.$
Second, using that $j\geqslant a(x)$ and $a(x)\to\infty$ as $x\to\infty$, as well as Theorem 3.2 of [gantert2018large], we have for some $c_{3},c_{4}>0$ large and independent of $x$, such that for all $x$ large enough,
$\hskip 14.22636pt\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x},\penalty 10000\ \bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)^{c}\}\right)$
@@
$\leqslant d\,\mathbb{P}\left(\max_{v\in V_{j}}|\eta^{(1)}_{v,j}(j)|\geqslant\frac{(c_{3}-1)j}{\sqrt{d}}\right)$
@@
For any $\varepsilon>0$ and $a(x)\to\infty$, the right-hand side is $\varepsilon$ for $x$ large enough. This completes the proof of the lower bound (24).
@@ 2.3 Tightness of first passage times: proof of Theorem 2
We build upon ideas from [mcdiarmid1995minimal], while in contrast, we do not assume uniform boundedness of the increments. In the following, we write $C_{1},C_{2},C_{3},\dots$ as large constants and $c_{3},c_{4},\dots$ as small constants that may depend on the law of $\bm{\xi}$} and the branching rate. [^8] Not to be confused with the fixed constants $c_{1},c_{2}$ defined in Section 1.1.1. Here, we do not assume that $\bm{\xi}$} is spherically symmetric. We start with a few preparatory lemmas.
@@
Assume (A6). For any $C_{1}>0$ and $c_{3}>0$, there exists $C_{2}>0$ such that the following holds for $k$ large enough: let $(X_{j},\mathbf{Y}_{j})$ be a sequence of i.i.d. random variables with the same distribution as $\bm{\xi}$}, $C>C_{2}$, and $(F_{Ck},\mathbf{J}_{Ck})=\sum_{j=1}^{Ck}(X_{j},\mathbf{Y}_{j})$. Then uniformly for $k$ large enough and ${\mathbf{z}}\in\mathbb{R}^{d}$ such that $\left\lVert{\mathbf{z}}\right\rVert\leqslant C_{1}k$,
$\mathbb{P}((F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}})\gg e^{-c_{3}k}.$
@@
$\mathbb{P}(\#\{v\in V_{n}:\left\lVert\bm{\eta}_{v,n}(n)\right\rVert\leqslant c_{5}n\}<e^{c_{6}n}\mid S)\ll e^{-c_{7}n}.$ (27)
Indeed, given $c_{5}$, let $c_{8}=c_{5}/(6d\bar{c}_{1})$ where here $\bar{c}_{1}$ stands for the maximum linear speed of the BRW across all $d$ directions, so that by Theorem 3.2 of [gantert2018large], [^9] Theorem 3.2 of [gantert2018large] is stated in the Schröder case, where $p_{0}+p_{1}>0$. However, the proof for the upper bound of the large deviation probability does not rely on this assumption; see Remark 3.3(b) therein.
$\mathbb{P}\left(\exists v\in V_{c_{8}n},\penalty 10000\ \left\lVert\bm{\eta}_{v,c_{8}n}(c_{8}n)\right\rVert>\frac{c_{5}n}{2}\right)\ll e^{-c_{7}n}.$ (28)
Conditioning upon the survival event $S$, with exponentially small probability we have that $V_{c_{8}n}\leqslant e^{c_{9}n}$. Evidently, each $v\in V_{c_{8}n}$ will lead independently with probability >1/2>1/2 to a particle $w\in V_{n}$ where $\left\lVert\bm{\eta}_{w,n}(n)-\bm{\eta}_{w,n}(c_{8}n)\right\rVert\leqslant c_{5}n/2$. This along with standard binomial estimates proves (27).
We return to the proof. We apply first (27) with $n=k$ and $c_{5}=1$. Outside a set of exponentially small probability, we may consider a subcollection of $e^{c_{6}k}$ particles in $\{v\in V_{k}:\left\lVert\bm{\eta}_{v,k}(k)\right\rVert\leqslant k\}$, which we label by $v_{j},\penalty 10000\ 1\leqslant j\leqslant e^{c_{6}k}$. We next evolve the particles $v_{j}$ independently and check if they reach $k$ in the $x$-direction. It suffices to consider the presence of the particles evolved from $v_{j},\penalty 10000\ 1\leqslant j\leqslant e^{c_{6}k}$ in a ball of radius one and at most $k+k+1\leqslant 3k$ apart (from the locations of $v_{j}$). By Lemma 11 applied with $C_{1}=3$ and $c_{3}=c_{6}/2$, for each $j$ we have for some $C_{2}>0$, with $w_{j}$ denoting the location of (a fixed descendant of) $v_{j}$ after time $C_{2}k$, that
@@
Note that such a fixed path exists for at least half of $v_{j}$’s with overwhelming probability. In conclusion, we evolved at least $e^{c_{6}k}/2$ many particles independently at time $k$ for time $C_{2}k$, each resulting in a probability of $\gg e^{-c_{6}k/2}$ to land in $k,\infty)\times\mathbb{R}^{d-1}$. A standard binomial estimate yields that with overwhelming probability, there exist $e^{c_{6}k/4}$ particles at time $(C_{2}+1)k$ present in $k,\infty)\times\mathbb{R}^{d-1}$ (conditional on survival). This completes the proof. ∎
@@
In the proof of Theorem 2, we wish to use Lemma 13 to bound probabilities of the form $\mathbb{P}(\tau_{x}\geqslant\alpha,\penalty 10000\ \tau_{y}\leqslant\beta)$, where $\beta<\alpha$. This is intuitively possible since we may run the particle $v_{y}$ that reaches $B_{y}$ and bound the probability that it never reaches $B_{x}$ in a time longer than $\alpha-\beta$, where we have used the strong Markov property of the random walk. There is, however, a subtle issue for the survival of the branching process initiated by the particle $v_{y}$, as this event is almost independent of the survival event $S$ we are conditioning on (in the proof of Theorem 2). We will bypass this difficulty by excluding this termination event a priori, or by first having an exponential number $e^{cn}$ of particles in $B_{y}$ instead of a single particle $v_{y}$.
@@
Let $k=y/C_{3}$ and set $C_{5}\gg 1+C_{3}+C_{6}C_{7}$, all to be determined. Denote by $\tau^{(c_{11})}_{x+k}$ the waiting time until $e^{c_{11}k}$ many particles reach the ball $B_{x+k}$. Now note that
$\begin{split}\mathbb{P}(\tau_{x}>s-1)&\leqslant\mathbb{P}(s-1<\tau_{x}<s+C_{5}k)+\mathbb{P}\left(\tau^{(c_{11})}_{x+k}>s+C_{3}k+1+C_{6}C_{7}k\right)\\ &\hskip 71.13188pt+\mathbb{P}\left(\tau_{x}\geqslant s+C_{5}k,\penalty 10000\ \tau^{(c_{11})}_{x+k}\leqslant s+C_{3}k+1+C_{6}C_{7}k\right).\end{split}$ (31)
@@
$\mathbb{P}(\#\{v\in V_{C_{3}k}:\eta^{(1)}_{v,C_{3}k}(C_{3}k)\geqslant k\}<e^{c_{4}k})\leqslant q+o(e^{-c_{4}k}).$